let P(x) denote the statement "x-3>5" . What are the truth values? P(2) P(8) P(9)


1
Expert's answer
2021-05-06T16:13:28-0400

Let "P(x)" denote the statement " "x-3>5" " . Since "2-3=-1<5", we conclude that "P(2)" is false. Taking into account that "8-3=5", we conclude that "P(8)" is false. Since "9-3=6>5", we conclude that "P(9)" is true.